solve the following differential equation: 2x^2y(d^2y/dx^2)+4y^2=x^2(dy/dx)^2+2xy(dy/dx)

1
Expert's answer
2021-12-31T11:34:33-0500

"2x^2yy''+4y^2=x^2(y')^2+2xyy'"


"y=x^2z^2"

then:

"y=x^2z^2"

"y'=2xz^2+2x^2zz'"

"y''=2z^2+4xzz'+4xzz'+2x^2((z')^2+zz'')"


"2x^2x^2z^2(2z^2+4xzz'+4xzz'+2x^2((z')^2+zz''))+4x^4z^4="

"=x^2(2xz^2+2x^2zz')^2+2xx^2z^2(2xz^2+2x^2zz')"


"x^2z^2(z^2+4xzz'+x^2((z')^2+zz''))+x^2z^4="

"=x^2z^4+2x^3z^3z'+x^4z^2(z')^2+xz^2(xz^2+x^2zz')"


"z^2+4xzz'+x^2((z')^2+zz'')+z^2=z^2+2xzz'+x^2(z')^2+z^2+xzz'"

"xzz'+x^2zz''=0"

"xz(z'+xz'')=0"


"z'+xz''=0"

"z'=u"

"u+xu'=0"

"du\/u=-dx\/x"

"lnu=-lnx+lnc_1"

"u=c_1\/x"

"dz=c_1dx\/x"

"z=c_1lnx+c_2"


"y(x)=x^2(c_1lnx+c_2)^2"